You would like to buy a car in 2 years. The car will cost $10,000 and the bank offers semiannual compounding on your savings account at an annual interest rate of 3%. How much do you need to contribute to the account at the end of each six-month period?

answer: $2,444

Plz show working progress.
Answer #1

Solution:

Semiannual interest rate = 3%/2 = 1.50%

Let semiannual contribution = X

Nos of semiannual period = 2*2 = 4

Required contribution = $10,000 / Cumulative FV factor at 1.5% for 4 periods

= $10,000 / 1.061364 = $10,000 / 4.0909 = $2,444